# Christoph Bandt

### v1

@@ 5 More examples
Figure 7: Sierpinski tetrahedron and its automaton, $0\leq i<j\leq 3$
Sierpinski gasket and tetrahedron.
The digits are $0,1,...,n$ for the $n$-dimensional version. For $n=1$ we have Example 1. There are states $ij$ for all pairs of digits $i\not=j.$ Thus in the plane we have 6 states and in $n$ dimensions we have $n(n+1).$ There are only edges from $o$ to all other states and loops, exactly as in Example 1. See Figure 7.
@@
The $2\times 2$ square is the basis of commonly used ‘quadtree’ methods for image coding and processing. We use digits $0,1,2,3$ as indicated in Figure 8. There are four states $N,E,S,W$ for the four sides of the square and four states $NE,SE,SW,NW$ for the vertices. This can be considered as product of Example 1 with itself. A $k\times k$ square has $k^{2}$ digits, but the same states and edges. Cubes are also easy to construct.
@@
This class of examples is obtained from the $k\times k$ square by taking less than $k^{2}$ digits. The classical example is the Sierpinski carpet, with a hole in the middle of a $3\times 3$ square. There are a lot of recent papers on fractal squares, mainly by Chinese authors. See [14, 23, 33, 40, 41, 48, 49, 50] and their references. The automaton for fractal squares is the same as for the corresponding square, with less edge labels. Some states can disappear, as in Figure 9. The case of only two states was studied by Zhu and Yang [51] with automata similar to ours.
Figure 9 also shows an example of a ‘fractal triangle’. Every second triangle is put upside down together with its sides. The automaton for this example needs three states associated to the three sides of a triangle, and has two loops at each state. Related examples will need six states when triangles meet at their vertices.
@@
This is the automata-free formulation of self-similarity. In the setting of Proposition 6, it implies the equation (2) for the homeomorphisms $h_{j}(x)=\varphi^{-1}\tau_{j}\varphi(x),j=0,...,m-1$ from $X$ into $X.$ The address map and space $X$ are called post-critically finite if there are finitely many double addresses $(s,t)$ with $s_{1}\not=t_{1},$ and all these sequences are preperiodic. This holds true for Figures 14 and the Sierpinski tetrahedron, but not for the square, Example 2 and Figure 9. The following result is related to [9, Proposition 6] and [39].
Proposition 7
@@
Proof. If $\varphi$ is postcritically finite, each double address $(s,t)$ with $s_{1}\not=t_{1}$ can be written as $s=s_{1}...s_{n}\overline{u}$ and $t=t_{1}...t_{n}\overline{v}$ with a common $n$ and words $u=u_{1}...u_{k},v=v_{1}...v_{k}$ of equal length. Consider a path of directed edges $e_{j},j=1...n+k$ with labels $(s_{j},t_{j})$ for $j=1,...,n$ and $(u_{i},v_{i})$ for $j=n+i,i=1,...,k.$ Let the terminal point of the last edge be the initial point of $e_{n+1}.$ Doing this for all double addresses with $s_{1}\not=t_{1}$ and joining the paths at their initial point which we call $o,$ we get the automaton.
Now assume the automaton is given and there is no path between two directed cycles. In particular, cycles are disjoint. Then from $o$ to any cycle there are only finitely many paths of edges. Thus each cycle accepts only finitely many pairs of preperiodic addresses. Since each infinite path from $o$ must go through a cycle, this implies that $\varphi$ is postcritically finite. $\Box$
@@ 6 Automata for multiple addresses
Definition 1 does not require that all pairs of addresses $(s,t)$ of one point $x$ in $X$ are accepted by the automaton. Such pairs can also come from transitivity of the equivalence relation, say when $(s,s^{\prime})$ and $(t,s^{\prime})$ are accepted by the automaton for some sequence $s^{\prime}.$ In Figure 4 we added one more state and edge for completeness. However, this was not necessary. The incomplete automaton provides the correct quotient space $X.$
@@
Figure 10: In Figures 4 and 5 we actually have only triple addresses and no proper double addresses. Here are the automata $G_{3}$ for the triple addresses which can be constructed by the method in Section 6. Label $iii$ is abbreviated as $i$.
@@
Usually incomplete automata are much simpler than the complete version, as shown by Figure 13 below compared to Figure 8 above. And given an automaton, we do not know whether it is complete or incomplete. We now describe an abstract algorithm which determines automata for triple and multiple addresses from a given automaton $G.$ It decides completeness of $G$ and also finds cases where we have no proper double addresses, only multiple addresses, as in Figure 10. In some way, we unfold the automaton $G,$ generalizing Gilbert’s work on triple addresses of complex number systems with base $-n+i$ [19].
Theorem 2 (Automata for triple and multiple addresses)
@@
Let $G=G_{2}$ be a topology-generating automaton with bounded equivalence classes and $\varphi:S\to X$ the corresponding address map. All multiple addresses of points in $X$ can be determined by automata $G_{3},G_{4},...$ derived from $G,$ without any geometric knowledge of $X.$ The automaton $G_{k}$ will accept all $k$-tuples of equivalent addresses which are not contained in a $(k+1)$-tuple of equivalent addresses, and also some others.
@@
Proof. For triple addresses, we have a product construction: $G_{3}$ will be a subset of $G_{2}\times G_{2}=(V^{2},D^{3},E_{3},oo)$ where the states are denoted $bc$ with $b,c\in V.$ There is an edge $(bc,b^{\prime}c^{\prime})\in E_{3}$ with label $ijk$ if $(b,b^{\prime})$ and $(c,c^{\prime})$ are edges in $G_{2}$ with labels $(i,j)$ and $(j,k),$ respectively. Two linked pairs of digits are coupled at their common item, which for paths will then realize the transitivity of the equivalence relation. To obtain $G_{3},$ we omit all vertices from $G_{2}\times G_{2}$ without outgoing edge and those components of the graph on which there are no edge paths with three different addresses. This leads from Figures 4 and 5 to Figure 10, for instance.
@@
Now we describe the construction of $G_{4},G_{5}$ etc. The algorithm for extending $G_{n}$ to $G_{n+1}$ is based on taking products $G_{n}\times G_{2}.$ That is, we join states $c$ of $G_{2}$ to $(n-1)$-tuples $b=b_{1}...b_{n-1}$ which describe states of $G_{n}.$ We draw edges when the last coordinate $j$ of the label between $b$ and $b^{\prime}$ agrees with the first coordinate of the label $(j,k)$ between $c$ and $c^{\prime}.$ In this way we extend the $n$ digits at certain edges of $G_{n}$ by another digit. If this works for an infinite path of edges, we have extended the number of addresses from $n$ to $n+1.$ If no further address can be joined to an equivalence class, it is complete and belongs to $G_{n}.$ If we can add an address, we get a class which belongs to $G_{n+1}.$ Since the size of equivalence classes is assumed to be bounded, the algorithm stops after finite time.
@@
Two problems must be mentioned. On the one hand, the new address which we find may often coincide with a previous address. The cleaning of the graph needs lexikographic ordering and recursive procedures which we do not discuss. On the other hand, it is not enough to take only the last entry of the labels of $G_{n}$ as $j.$ For the examples below, a new address has sometimes to be appended at the first entry. In general, we cannot assume that transitivity is realized in a chain. We must check the product with $G_{2}$ for each fixed coupling coordinate $j$ between 1 and $n,$ and then take the union of the $n$ graphs with their initial states identified. This is possible since in any instance we add at most one new address. In the general case, there is no meaning in writing the states of $G_{n}$ as tuples of states in $G_{2}.$ We can take any other names. There is another simplification in extending $G_{n}$ at each coordinate $j.$ We need only one permutation of each set of addresses in $G_{n},$ because $G_{2}$ satisfies property 3 in definition 1. Thus the graphs $G_{k}$ with $k\geq 3$ are represented in asymmetric form, simpler than $G_{2},$ as can be seen in Figure 10. Only when a sequence of edges leads back to a state with permutation of the addresses, we have to replicate the state.
@@
Any address which is equivalent to one address in an $n$-tuple of addresses in $G_{n}$ will be found in this way. And if such a new address does not exist, the $n$-tuple is confirmed as being complete and belonging to $G_{n}.$ Every incomplete address will be extended and eventually be completed, since $n$ is bounded by assumption. This concludes the proof. $\Box$
@@
The method must be augmented by ordering and graph cleaning procedures for efficient programming. In the present form, it works well for small examples, as Figure 13 below and the Example 4 in Section 9. We now discuss a more complicated case.
Example 3 (An incomplete automaton generating a triangle)
@@
Figure 11 shows an automaton with three states and three digits. The corresponding space $X$ is a well-known triangle. It is a self-similar set and the fundamental domain of the Coxeter group $\tilde{G_{2}}$ generated by the reflections at the three sides of the triangle. The neighbor graph mentioned in Section 9 is a complete automaton describing all pairs of equivalent addresses. For this example it has 16 states and 42 arrows which explains why we prefer the incomplete automaton. Some multiple addresses can be determined by eyesight from the geometry of triangle pieces. Note that two vertices of the triangle have addresses $\overline{0}$ and $\overline{2}$ and all multiple addresses will end with one of these suffixes. There are points with 4, 6 and even 12 addresses which correspond to the vertices of the corresponding crystallographic tiling.
@@
Figure 11: An incomplete automaton generating a self-similar triangle. The states of the automaton correspond to the three sides $a,b,c.$ In the small triangle $0c$ is identified with $1c$ and $1a$ with $2a.$ The big triangle shows third-level pieces around points with 4, 6 and 12 addresses.
@@
We construct $G_{3}$ for Example 3 by taking $G_{2}\times G_{2}.$ The states $ac$ and $ca$ have no outgoing edges. Moreover, the states $oo,aa,bb,cc$ are connected by edges with first label equal to the third. Since $aa,bb,cc$ do not lead to other states, they can be omitted. Thus from 16 states we are left with $oo$ and 10 other states of $G_{3}.$ Because of symmetry, we do not show the states $oa,ob,oc$ in Figure 12. Labels $iii$ are abbreviated $i.$
@@
We have paths of two edges from $oo$ to $bc$ and $ba$ which directly yield triple addresses. Paths through $ao$ yield $10\overline{2}\sim 20\overline{2}\sim 21\overline{2}$ and $20\overline{2}\sim 10\overline{2}\sim 11\overline{2}$ for the two labels of the edge $(oo,ao).$ Similarly, the paths from $oo$ over $co$ to $bc$ yield $01\overline{2}\sim 11\overline{2}\sim 10\overline{2}$ and $11\overline{2}\sim 01\overline{2}\sim 00\overline{2}.$ For the paths to $ba$ two pairs of labels can be combined: $01\overline{0}\sim 11\overline{0}\sim 12\overline{0}$ and $02\overline{0}\sim 12\overline{0}\sim 11\overline{0}$ and $11\overline{0}\sim 01\overline{0}\sim 02\overline{0}$ and finally $12\overline{0}\sim 02\overline{0}\sim 01\overline{0}$ since In Figure 11, these triples correspond to three consecutive gray third-level pieces at the point marked between pieces 0 and 1. The other four triple addresses correspond to consecutive blue triangles around the marked point on the basic line. We see that the triples are not complete: the paths to $ba$ correspond to a quadruple and the paths to $bc$ to a $6$-tuple of addresses. Thus we do not need $G_{3},$ but $G_{4}$ and $G_{6}.$
@@
Figure 12: Automata $G_{3}$ and $G_{4}$ for example 3. Ignoring loops at $oo$ and paths through $bo,$ there are eight triple addresses which correspond to the common endpoints of three consecutive third-level triangles in Figure 11. These triples can all be extended to quadrupels. The quadrupel addresses leading to $aba$ in $G_{4}$ are complete, those leading to $cbc$ can be further extended.
We construct $G_{4}$ starting with the edge from $oo$ to $a0$ in $G_{3}.$ For 122 a new item can be appended in front, but not at the end. However, 211 can be extended to both sides, yielding edges from $ooo$ to $aoc$ and $oao.$ The first edge determines a single path while the second one determines the labelling of the rest of the network, as shown on the right of Figure 12. All other couplings of addresses, for instance appending to 122 or taking the other label of edges to and from $c0$ in $G_{3},$ will lead to repetitions, up to permutation of the addresses.
All triple addresses have been extended, so $G_{3}$ is now irrelevant. In $G_{4}$ in Figure 12 the paths leading to $aba$ describe quadruples which are complete. There are two possible digits for the starting edge at $ooo$ and two possible digits for the final edge at $aba,$ so there can be only four addresses. When we try to append another address, it will be a repetition. Thus the part of $G_{4}$ which leads to $aba,$ comprising six vertices, is the graph of complete quadruple addresses.
@@
$G_{6}:\quad\textstyle ooooo{221100\atop\longrightarrow}oaoco{100110\atop\longrightarrow}cbcbc{2\atop\longleftrightarrow}bcbcb\,.$
This path describes the six addresses of the marked point on the basic line, and, together with the loops at the initial state, all points in $X$ with six and more addresses. Thus this is already the graph $G_{6}.$ The points which lie on the sides of the triangle $X$ have exactly six addresses, the other have twelve. The sides of type $b$ on the boundary are $b,0b,1b,2b.$ We have the self-similarity equation $b=00b\cup 21b\cup 22b.$ So the boundary has an automatic structure, with address language $\{00,21,22\}^{*}\{\lambda,0,1,2\}$ where $\lambda$ is the empty word. Thus we can replace $ooooo$ in $G_{6}$ by a small automaton with five states to describe exactly the complete six-tuple addresses of points on the boundary. This is not contained in our general method which also produces a larger version of $G_{6}.$
To determine the twelve-tuple addresses, we really need the states with several $o$ as in Figure 12. We have to amalgamate two six-tuple addresses along the line $1a=2a$ (marked point in Figure 11) or along $0c=1c$ (two points on both sides of the marked point). Here is one version of the automaton, with notation $v^{n}=vv...v.$ All twelve-tuples are complete.
$G_{12}:\quad\textstyle o^{11}\ \ {1^{6}2^{6}\atop\longrightarrow}$ $o^{5}ao^{5}$ $\textstyle{0\atop\longleftrightarrow}\ o^{5}bo^{5}\ {(221100)^{2}\atop\longrightarrow}oaocoaocoao{(1001)^{3}\atop\longrightarrow}cbcbcbcbcbc{2\atop\longleftrightarrow}bcbcbcbcbcb\,.$
@@
This completes the discussion of Example 3. The simpler case of the square is illustrated in Figure 13. Example 4 with Figure 14 leads to $G_{3},G_{4},$ and $G_{6}.$
Figure 13: Top: a simple incomplete automaton for the $2\times 2$ square and the corresponding automaton of 4-tuple addresses. Bottom: the finite topological spaces $X_{1}$ and $X_{2}$ approximating the square $X.$ Gray circles are open points corresponding to $D$ and $D^{2}.$ Black circles are closed points representing the 4-tuple addresses, and line segments correspond to double addresses. The automata organize this well-known pattern [27] in a hierarchical way.

### v2

@@ 5 More examples
Figure 7: Sierpiński tetrahedron and its automaton, $0\leq i<j\leq 3$.
Sierpiński gasket and tetrahedron.
The digits are $0,1,\dots,n$ for the $n$-dimensional version. For $n=1$ we have Example 2.1. There are states $ij$ for all pairs of digits $i\not=j$. Thus in the plane we have 6 states and in $n$ dimensions we have $n(n+1)$. There are only edges from $o$ to all other states and loops, exactly as in Example 2.1. See Figure 7.
@@
The $2\times 2$ square is the basis of commonly used ‘quadtree’ methods for image coding and processing. We use digits $0,1,2,3$ as indicated in Figure 8. There are four states $N,E,S,W$ for the four sides of the square and four states $NE,SE,SW,NW$ for the vertices. This can be considered as a product of Example 2.1 with itself. A $k\times k$ square has $k^{2}$ digits, but the same states and edges. Cubes are also easy to construct.
@@
This class of examples is obtained from the $k\times k$ square by taking less than $k^{2}$ digits. The classical example is the Sierpiński carpet, with a hole in the middle of a $3\times 3$ square. There are a lot of recent papers on fractal squares, mainly by Chinese authors. See [15, 24, 34, 42, 43, 50, 51, 52] and their references. The automaton for fractal squares is the same as for the corresponding square, with fewer edge labels. Some states can disappear, as in Figure 9. The case of only two states was studied by Zhu and Yang [53] with automata similar to ours.
Figure 9 also shows an example of a ‘fractal triangle’. Every second triangle is put upside down together with its sides. The automaton for this example needs three states associated with the three sides of a triangle, and has two loops at each state. Related examples will need six states when triangles meet at their vertices.
@@
This is the automata-free formulation of self-similarity. In the setting of Proposition 4.1, it implies the equation (2) for the homeomorphisms $h_{j}(x)=\varphi^{-1}\tau_{j}\varphi(x),j=0,\dots,m-1$ from $X$ into $X$. The address map and space $X$ are called post-critically finite if there are finitely many double addresses $(s,t)$ with $s_{1}\not=t_{1}$, and all these sequences are preperiodic. This holds for Figures 14 and the Sierpiński tetrahedron, but not for the square, Example 3.4 and Figure 9. The following result is related to [10, Proposition 6] and [41].
Proposition 5.1.
@@
Proof 5.2.
If $\varphi$ is postcritically finite, each double address $(s,t)$ with $s_{1}\not=t_{1}$ can be written as $s=s_{1}\dots s_{n}\overline{u}$ and $t=t_{1}\dots t_{n}\overline{v}$ with a common $n$ and words $u=u_{1}\dots u_{k},v=v_{1}\dots v_{k}$ of equal length. Consider a path of directed edges $e_{j},j=1\dots n+k$ with labels $(s_{j},t_{j})$ for $j=1,\dots,n$ and $(u_{i},v_{i})$ for $j=n+i,i=1,\dots,k$. Let the terminal point of the last edge be the initial point of $e_{n+1}$. Doing this for all double addresses with $s_{1}\not=t_{1}$ and joining the paths at their initial point which we call $o$, we get the automaton.
Now assume the automaton is given and there is no path between two directed cycles. In particular, cycles are disjoint. Then from $o$ to any cycle there are only finitely many paths of edges. Thus each cycle accepts only finitely many pairs of preperiodic addresses. Since each infinite path from $o$ must go through a cycle, this implies that $\varphi$ is postcritically finite.
@@ 6 Automata for multiple addresses
Definition 2.4 does not require that all pairs of addresses $(s,t)$ of one point $x$ in $X$ are accepted by the automaton. Such pairs can also come from the transitivity of the equivalence relation, say when $(s,s^{\prime})$ and $(t,s^{\prime})$ are accepted by the automaton for some sequence $s^{\prime}$. In Figure 4 we added one more state and edge for completeness. However, this was not necessary. The incomplete automaton provides the correct quotient space $X$.
@@
Figure 10: In Figures 4 and 5 we have only triple addresses and no proper double addresses. Here are the automata $G_{3}$ for the triple addresses which can be constructed by the method in Section 6. Label $iii$ is abbreviated as $i$.
@@
Usually, incomplete automata are much simpler than the complete version, as shown by Figure 13 below compared to Figure 8 above. And given an automaton, we do not know whether it is complete or incomplete. We now describe an abstract algorithm that determines automata for triple and multiple addresses from a given automaton $G$. It decides the completeness of $G$ and also finds cases where we have no proper double addresses, only multiple addresses, as in Figure 10. In some way, we unfold the automaton $G$, generalizing Gilbert’s work on triple addresses of complex number systems with base $-n+i$ [20].
Theorem 6.1 (Automata for triple and multiple addresses).
@@
Let $G=G_{2}$ be a topology-generating automaton with finite equivalence classes, with size smaller than a constant $C$. Let $\varphi:S\to X$ be the corresponding address map. All multiple addresses of points in $X$ can be determined by automata $G_{3},G_{4},\dots$ derived from $G$, without any geometric knowledge of $X$. The automaton $G_{k}$ will accept all $k$-tuples of equivalent addresses that are not contained in a $(k+1)$-tuple of equivalent addresses.
Proof 6.2.
@@
For triple addresses, we have a product construction: $G_{3}$ will be a subset of $G_{2}~{}\times~{}G_{2}=(V^{2},D^{3},E_{3},oo)$ where the states are denoted $bc$ with $b,c\in V$. There is an edge $(bc,b^{\prime}c^{\prime})\in E_{3}$ with label $ijk$ if $(b,b^{\prime})$ and $(c,c^{\prime})$ are edges in $G_{2}$ with labels $(i,j)$ and $(j,k)$, respectively. Two linked pairs of digits are coupled at their common item, which for paths will then realize the transitivity of the equivalence relation. To obtain $G_{3}$, we omit all vertices from $G_{2}\times G_{2}$ without outgoing edge and those components of the graph on which there are no edge paths with three different addresses. This leads from Figures 4 and 5 to Figure 10, for instance.
@@
Now we describe the construction of $G_{4},G_{5}$, etc. The algorithm for extending $G_{n}$ to $G_{n+1}$ is based on taking products $G_{n}\times G_{2}$. That is, we join states $c$ of $G_{2}$ to $(n-1)$-tuples $b=b_{1}\dotsb_{n-1}$ which describe states of $G_{n}$. We draw edges when the last coordinate $j$ of the label between $b$ and $b^{\prime}$ agrees with the first coordinate of the label $(j,k)$ between $c$ and $c^{\prime}$. In this way, we extend the $n$ digits at certain edges of $G_{n}$ by another digit. If this works for an infinite path of edges, we have extended the number of addresses from $n$ to $n+1$. If no further address can be joined to an equivalence class, it is complete and belongs to $G_{n}$. If we can add an address, we get a class which belongs to $G_{n+1}$. Since the size of equivalence classes is assumed to be bounded, the algorithm stops after finite time.
@@
Two problems must be mentioned. On the one hand, the new address which we find may often coincide with a previous address. The cleaning of the graph needs lexicographic ordering and recursive procedures which we do not discuss. On the other hand, it is not enough to take only the last entry of the labels of $G_{n}$ as $j$. For the examples below, a new address has sometimes to be appended at the first entry. In general, we cannot assume that transitivity is realized in a chain. We must check the product with $G_{2}$ for each fixed coupling coordinate $j$ between 1 and $n$, and then take the union of the $n$ graphs with their initial states identified. This is possible since in any instance we add at most one new address. In the general case, there is no meaning in writing the states of $G_{n}$ as tuples of states in $G_{2}$. We can take any other names. There is another simplification in extending $G_{n}$ at each coordinate $j$. We need only one permutation of each set of addresses in $G_{n}$, because $G_{2}$ satisfies property 3 in Definition 2.4. Thus the graphs $G_{k}$ with $k\geq 3$ are represented in asymmetric form, simpler than $G_{2}$, as can be seen in Figure 10. Only when a sequence of edges leads back to a state with a permutation of the addresses, we have to replicate the state.
@@
Any address that is equivalent to one address in an $n$-tuple of addresses in $G_{n}$ will be found in this way. And if such a new address does not exist, the $n$-tuple is confirmed as being complete and belonging to $G_{n}$. Every incomplete address will be extended and eventually be completed since $n$ is bounded by assumption. This concludes the proof.
@@
The method must be augmented by ordering and graph cleaning procedures for efficient programming. In the present form, it works well for small examples, as in Figure 13 below and Example 9.5 in Section 9. We now discuss a more complicated case.
Example 6.3 (An incomplete automaton generating a triangle).
@@
Figure 11 shows an automaton with three states and three digits. The corresponding space $X$ is a well-known triangle. It is a self-similar set and the fundamental domain of the Coxeter group $\tilde{G_{2}}$ generated by the reflections at the three sides of the triangle. The neighbor graph mentioned in Section 9 is a complete automaton describing all pairs of equivalent addresses. For this example it has 16 states and 42 arrows which explains why we prefer the incomplete automaton.d Some multiple addresses can be determined by eyesight from the geometry of triangle pieces. Note that two vertices of the triangle have addresses $\overline{0}$ and $\overline{2}$ and all multiple addresses will end with one of these suffixes. There are points with 4, 6, and even 12 addresses which correspond to the vertices of the corresponding crystallographic tiling.
@@
Figure 11: An incomplete automaton generating a self-similar triangle. The states of the automaton correspond to the three sides $a,b,c$. In the small triangle $0c$ is identified with $1c$ and $1a$ with $2a$. The big triangle shows third-level pieces around points with 4, 6, and 12 addresses.
@@
We construct $G_{3}$ for Example 6.3 by taking $G_{2}\times G_{2}$. The states $ac$ and $ca$ have no outgoing edges. Moreover, the states $oo,aa,bb,cc$ are connected by edges with the first label equal to the third. Since $aa,bb,cc$ do not lead to other states, they can be omitted. Thus from 16 states, we are left with $oo$ and 10 other states of $G_{3}$. Because of symmetry, we do not show the states $oa,ob,oc$ in Figure 12. Labels $iii$ are abbreviated $i$.
@@
We have paths of two edges from $oo$ to $bc$ and $ba$ which directly yield triple addresses. Paths through $ao$ yield $10\overline{2}\sim 20\overline{2}\sim 21\overline{2}$ and $20\overline{2}\sim 10\overline{2}\sim 11\overline{2}$ for the two labels of the edge $(oo,ao)$. Similarly, the paths from $oo$ over $co$ to $bc$ yield $01\overline{2}\sim 11\overline{2}\sim 10\overline{2}$ and $11\overline{2}\sim 01\overline{2}\sim 00\overline{2}$. For the paths to $ba$ two pairs of labels can be combined: $01\overline{0}\sim 11\overline{0}\sim 12\overline{0}$ and $02\overline{0}\sim 12\overline{0}\sim 11\overline{0}$ and $11\overline{0}\sim 01\overline{0}\sim 02\overline{0}$ and finally $12\overline{0}\sim 02\overline{0}\sim 01\overline{0}$. In Figure 11, these triples correspond to three consecutive gray third-level pieces at the point marked between pieces 0 and 1. The other four triple addresses correspond to consecutive blue triangles around the marked point on the basic line. We see that the triples are not complete: the paths to $ba$ correspond to a quadruple and the paths to $bc$ to a $6$-tuple of addresses. Thus we do not need $G_{3}$, but $G_{4}$ and $G_{6}$.
@@
Figure 12: Automata $G_{3}$ and $G_{4}$ for example 6.3. Ignoring loops at $oo$ and paths through $bo$, there are eight triple addresses that correspond to the common endpoints of three consecutive third-level triangles in Figure 11. These triples can all be extended to quadruples. The quadruple addresses leading to $aba$ in $G_{4}$ are complete, those leading to $cbc$ can be further extended.
We construct $G_{4}$ starting with the edge from $oo$ to $a0$ in $G_{3}$. For 122 a new item can be appended in front, but not at the end. However, 211 can be extended to both sides, yielding edges from $ooo$ to $aoc$ and $oao$. The first edge determines a single path while the second one determines the labeling of the rest of the network, as shown on the right of Figure 12. All other couplings of addresses, for instance appending to 122 or taking the other label of edges to and from $c0$ in $G_{3}$, will lead to repetitions, up to permutation of the addresses.
All triple addresses have been extended, so $G_{3}$ is now irrelevant. In $G_{4}$ in Figure 12 the paths leading to $aba$ describe quadruples that are complete. There are two possible digits for the starting edge at $ooo$ and two possible digits for the final edge at $aba$, so there can be only four addresses. When we try to append another address, it will be a repetition. Thus the part of $G_{4}$ which leads to $aba$, comprising six vertices, is the graph of complete quadruple addresses.
@@
This path describes the six addresses of the marked point on the basic line, and, together with the loops at the initial state, all points in $X$ with six and more addresses. Thus this is already the graph $G_{6}$. The points that lie on the sides of the triangle $X$ have exactly six addresses, the other have twelve. The sides of type $b$ on the boundary are $b,0b,1b,2b$. We have the self-similarity equation $b=00b\cup 21b\cup 22b$. So the boundary has an automatic structure, with address language $\{00,21,22\}^{*}\{\lambda,0,1,2\}$ where $\lambda$ is the empty word. Thus we can replace $ooooo$ in $G_{6}$ by a small automaton with five states to describe exactly the complete six-tuple addresses of points on the boundary. This is not contained in our general method which also produces a larger version of $G_{6}$.
To determine the twelve-tuple addresses, we really need the states with several $o$ as in Figure 12. We have to amalgamate two six-tuple addresses along the line $1a=2a$ (marked point in Figure 11) or along $0c=1c$ (two points on both sides of the marked point). Here is one version of the automaton, with notation $v^{n}=vv\dots v$. All twelve tuples are complete.
[…]
This completes the discussion of Example 6.3. The simpler case of the square is illustrated in Figure 13. Example 9.5 with Figure 14 leads to $G_{3},G_{4}$, and $G_{6}$.
Figure 13: Top: a simple incomplete automaton for the $2\times 2$ square and the corresponding automaton of 4-tuple addresses. Bottom: the finite topological spaces $X_{1}$ and $X_{2}$ approximating the square $X$. Gray circles are open points corresponding to $D$ and $D^{2}$. Black circles are closed points representing the 4-tuple addresses, and line segments correspond to double addresses. The automata organize this pattern [28] in a hierarchical way.